OpenStudy (faman39):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $130. This year, he paid $140. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

%increase = difference/ original

OpenStudy (anonymous):

@Faman39 Please show me if you understand it?

OpenStudy (faman39):

Nope still couldnt understand :(

OpenStudy (baldymcgee6):

okay, what @Chlorophyll is saying is... %increase = (difference/ original) * 100 difference: the difference between the two prices. 140-130 = original: the original price paid for the plane ticket. $130 What part of this do you not understand?

OpenStudy (faman39):

after subtracting what do we need to do ?

OpenStudy (baldymcgee6):

We need to use our formula %increase = (difference/ original) * 100

OpenStudy (faman39):

130/10? is it like this?

OpenStudy (faman39):

I guess not, i m sorry guys

OpenStudy (faman39):

:(

OpenStudy (anonymous):

difference/ original . = (140-130 ) / 130

OpenStudy (baldymcgee6):

\[\LARGE \color{blue} {\%increase = \frac{ difference }{ original }}\]

OpenStudy (baldymcgee6):

Make sure you put the difference ontop. :)

OpenStudy (faman39):

7.6923?

OpenStudy (anonymous):

Then multiply the result for 100 to get the %increase :)

OpenStudy (anonymous):

@baldymcgee6 Thanks for making the formula colorful enjoyable to be @Faman39 memorized

OpenStudy (faman39):

7692.30?

OpenStudy (anonymous):

Sorry, 10 * 100 / 130 = 100/ 13 = 7.69%

OpenStudy (faman39):

Thank you guys :( sorry for giving so much trouble, i m so sorry guys, i didnt want to trouble , but i m so dumbest person, thats y, i couldnt get anything